jason has invested a total of $25,000 in two accounts and at the end of one year earned $2,600 in simple intrest. one account earns 7% simple intrest and the other earns 12% simple intrest. What is the amount invested in each account?
Answer by checkley77(12569)

.12x+.07(25,000-x)=2,600
.12x+1,750-.07x=2,600
.05x=2,600-1,750
.05x=850
x=850/.05
x=17,000 invested @ 12%.
25,000-17,000=8,000 invested @ 7%.
proof:
.12*17,000+.07*8,000=2,600
2,040+560=2,600
2,600=2,600